# On polynomial interpolation in the monomial basis

### v4

@@
In this paper, we show that the monomial basis is generally as good as a well-conditioned polynomial basis for interpolation, provided that the condition number of the Vandermonde matrix is smaller than the reciprocal of machine epsilon.
@@ 1 Introduction
Function approximation has been a central topic in numerical analysis since its inception. One of the most effective methods for approximating a function $F:[-1,1]\to\mathbbm{R}$ is the use of an interpolating polynomial $P_{N}$ of degree $N$ which satisfies $P_{N}(x_{j})=F(x_{j})$ for a set of $(N+1)$ collocation points $\{x_{j}\}_{j=0,1,\dots,N}$. In practice, the collocation points are typically chosen to be the Chebyshev points, and the resulting interpolating polynomial, known as the Chebyshev interpolant, is a nearly optimal approximation to $F$ in the space of polynomials of degree at most $N$ [27]. A common basis for representing the interpolating polynomial $P_{N}$ is the Lagrange polynomial basis, and the evaluation of $P_{N}$ in this basis can be done stably using the Barycentric interpolation formula [8, 19]. Some other commonly used bases are Newton polynomials, Chebyshev polynomials, and Legendre polynomials. Alternatively, the monomial basis can be used to represent $P_{N}$, such that $P_{N}(x)=\sum_{k=0}^{N}a_{k}x^{k}$ for some coefficients $\{a_{k}\}_{k=0,1,\dots,N}$. The computation of the monomial coefficient vector $a:=(a_{0},a_{1},\dots,a_{N})^{T}\in\mathbbm{R}^{N+1}$ of the interpolating polynomial $P_{N}$ requires the solution to a linear system $Va=f$, where
$V:=\begin{pmatrix}1&x_{0}&x_{0}^{2}&\cdots&x_{0}^{N}\\ 1&x_{1}&x_{1}^{2}&\cdots&x_{1}^{N}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&x_{N}&x_{N}^{2}&\cdots&x_{N}^{N}\\ \end{pmatrix}\in\mathbbm{R}^{(N+1)\times(N+1)}$ (5)
@@
is a Vandermonde matrix, and $f:=\bigl{(}F(x_{0}),F(x_{1}),\dots,F(x_{N})\bigr{)}^{T}\in\mathbbm{R}^{N+1}$ is a vector of the function values of $F$ at the $(N+1)$ collocation points on the interval $[-1,1]$. It is well-known that, given any set of real collocation points, the condition number of a Vandermonde matrix grows at least exponentially as $N$ increases [7]. It follows that the numerical solution to this linear system is highly inaccurate when $N$ is not small, and, as a result, this algorithm for constructing $P_{N}$ is often considered to be unstable. But, is this really the case? Let $\{x_{j}\}_{j=0,1,\dots,N}$ be the set of $(N+1)$ Chebyshev points on the interval $[-1,1]$, and consider the case where $F(x)=\cos(2x+1)$. We solve the resulting Vandermonde system using LU factorization with partial pivoting. In Figure 0(a), we present a comparison between the approximation error of the computed monomial expansion (labeled as “Monomial”) and the approximation error of the Chebyshev interpolant evaluated using the Barycentric interpolation formula (labeled as “Lagrange”). One can observe that the computed monomial expansion is, surprisingly, as accurate as the Chebyshev interpolant evaluated using the Barycentric interpolation formula (which is accurate up to machine precision), despite the huge condition number of the Vandermonde matrix reported in Figure 0(b).
(a) $\cos(2x+1)$
(b) Condition number
@@
Figure 1: Polynomial interpolation of $\cos(2x+1)$ in the monomial basis. The $x$-axis label $N$ denotes the order of approximation. The $y$-axis label “Error” denotes the $L^{\infty}$ approximation error over $[-1,1]$, which is estimated by comparing the approximated function values at $10000$ equidistant points over $[-1,1]$ with the true function values.
@@
What happens when the function $F$ becomes more complicated? In Figure 2, we compare the accuracy of the two approximations when $F(x)=\cos(8x+1)$ and when $F(x)=\cos(12x+1)$. Initially, the computed monomial expansion is as accurate as the Chebyshev interpolant evaluated using the Barycentric interpolation formula. However, the convergence of polynomial interpolation in the monomial basis stagnates after reaching a certain error threshold. Furthermore, it appears that, the more complicated a function is, the larger that error threshold becomes. But what does it mean for a function to be complicated in this context? Consider the case where the function requires an even higher-order Chebyshev interpolant in order to be approximated to machine precision. In Figure 3, we compare the accuracy of the two approximations when $F(x)=\frac{1}{x-\sqrt{2}}$ and when $F(x)=\frac{1}{x-0.5i}$. These two functions each have a singularity in a neighborhood of the interval $[-1,1]$, and Chebyshev interpolants of degree $\geq 40$ are required to approximate them to machine precision. Yet, no stagnation of convergence is observed. In Figure 4, we consider the case where $F$ is a non-smooth function, and we find that the accuracy of the two approximations is, again, the same. Based on all of the previous examples, we conclude that polynomial interpolation in the monomial basis is not as unstable as it appears, and has some subtleties lurking around the corner that are worth further investigation.
(a) $\cos(8x+1)$
(b) $\cos(12x+1)$
@@
Figure 2: Polynomial interpolation of more complicated functions in the monomial basis.
(a) $\frac{1}{x-\sqrt{2}}$
(b) $\frac{1}{x-0.5i}$
@@
Figure 3: Polynomial interpolation of functions with a singularity near the interval $[-1,1]$ in the monomial basis.
(a) ${\lvert x+0.1\rvert}^{2.5}$
(b) ${\lvert\sin(5x)\rvert}^{3}$
@@
Figure 4: Polynomial interpolation of non-smooth functions in the monomial basis.
@@
These seemingly mysterious experiments can be explained partially from the point of view of backward error analysis. Indeed, the forward error ${\lVert a-\widehat{a}\rVert}_{2}$ of the numerical solution $\widehat{a}$ to the Vandermonde system $Va=f$ can be huge, but it is the backward error, i.e., ${\lVert V\widehat{a}-f\rVert}_{2}$, that matters for the accuracy of the approximation. This is because a small backward error implies that the difference between the computed monomial expansion, which we denote by $\widehat{P}_{N}$, and the exact interpolating polynomial, $P_{N}$, is a polynomial that approximately vanishes at all of the collocation points. When the Lebesgue constant associated with the collocation points is small (which is the case for the Chebyshev points), the polynomial $P_{N}-\widehat{P}_{N}$ is bounded uniformly by the backward error times a small constant. As a result, we bound the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}$ by the following inequality:
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}\leq{\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}+{\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}.$ (6)
@@
We refer to the first and the second terms on the right-hand side of (6) as the polynomial interpolation error and the backward error, respectively. When the backward error is smaller than the polynomial interpolation error, the monomial approximation error is dominated by the polynomial interpolation error, and the use of a monomial basis does not incur any additional loss of accuracy. Once the polynomial interpolation error becomes smaller than the backward error, the convergence of the approximation stagnates. For example, in Figure 2(a), we verify numerically that the backward error is around the size of machine epsilon for all $N\leq 43$, so stagnation is not observed, and polynomial interpolation in the monomial basis is as accurate as polynomial interpolation in the Lagrange basis, evaluated by the Barycentric interpolation formula. On the other hand, in Figure 1(a), the backward error is around the size of $10^{-13}$ for $N\geq 20$, which leads to stagnation once the polynomial interpolation error is less than $10^{-13}$.
@@
The explanation above brings up a new question: when will the backward error be small? When a backward stable linear system solver (e.g., LU factorization with partial pivoting) is used to solve the Vandermonde system $Va=f$, it is guaranteed that the numerical solution $\widehat{a}$ is the exact solution to the linear system
$(V+\delta V)\widehat{a}=f,$ (7)
for a matrix $\delta V\in\mathbbm{R}^{(N+1)\times(N+1)}$ that satisfies ${\lVert\delta V\rVert}_{2}\leq u\cdot\gamma$, where $u$ denotes machine epsilon and $\gamma=\mathcal{O}({\lVert V\rVert}_{2})$. It follows that the backward error, ${\lVert V\widehat{a}-f\rVert}_{2}$, of the numerical solution is bounded by $u\cdot\gamma{\lVert\widehat{a}\rVert}_{2}$. We note that $\gamma$ is typically small, so the backward error is essentially determined by the norm of the computed monomial coefficient vector. In fact, so long as $\kappa(V)\lesssim\frac{1}{u}$, one can show that the norm of the monomial coefficient vector computed by a backward stable solver is around the same size as the norm of the exact monomial coefficient vector of the interpolating polynomial. Therefore, in this case, the monomial approximation error can be quantified a priori using information about the interpolating polynomial, which implies that a theory of polynomial interpolation in the monomial basis can be developed.
@@
The rest of the paper is organized as follows. In Section 2, we analyze polynomial interpolation in the monomial basis over a smooth simple arc in the complex plane, with the interval as a special case, along with a number of numerical experiments. Our analysis shows that the monomial basis is similar to a well-conditioned polynomial basis for interpolation, provided that the condition number of the Vandermonde matrix is smaller than the reciprocal of machine epsilon. In Section 3, we present applications where the use of a monomial basis for interpolation offers a substantial advantage over other bases. In Section 4, we review related work, and discuss the generalization of our theory to higher dimensions.
@@ 2 Polynomial interpolation in the monomial basis
Let $\Gamma\subset\mathbbm{C}$ be a smooth simple arc, and let $F:\Gamma\to\mathbbm{C}$ be an arbitrary function. The $N$th degree interpolating polynomial, denoted by $P_{N}$, of the function $F$ for a given set of $(N+1)$ distinct collocation points $Z:=\{z_{j}\}_{j=0,1,\dots,N}\subset\Gamma$ can be expressed as $P_{N}(z)=\sum_{k=0}^{N}a_{k}z^{k}$, where the monomial coefficient vector $(a_{0},a_{1},\dots,a_{N})^{T}$ is the solution to the Vandermonde system
@@
For ease of notation, we denote the Vandermonde matrix by $V^{(N)}$, the monomial coefficient vector by $a^{(N)}$, and the corresponding right-hand side vector by $f^{(N)}$.
@@
${\lVert A^{-1}\rVert}_{2}\leq\frac{1}{\alpha\cdot{\lVert\delta A\rVert}_{2}},$ (21)
@@
The proof is complete by combining (25) and (26). $\blacksquare$
The following theorem provides upper bounds for the monomial approximation error. It can be viewed as a special case of frame approximation theory [3, 4]
@@
Let $\Gamma\subset\mathbbm{C}$ be a smooth simple arc, and let $F:\Gamma\to\mathbbm{C}$ be an arbitrary function. Suppose that $P_{N}$ is the $N$th degree interpolating polynomial of $F$ for a given set of $(N+1)$ distinct collocation points $Z:=\{z_{j}\}_{j=0,1,\dots,N}\subset\Gamma$. Clearly, the monomial coefficient vector $a^{(N)}$ of the polynomial $P_{N}$ is the solution to the Vandermonde system $V^{(N)}a^{(N)}=f^{(N)}$, where $V^{(N)}$ and $f^{(N)}$ have been previously defined in (20). Suppose further that there exists some constant $\gamma_{N}\geq 0$ such that the computed monomial coefficient vector $\widehat{a}^{(N)}=(\widehat{a}_{0},\widehat{a}_{1},\dots,\widehat{a}_{N})^{T}$ satisfies
$\bigl{(}V^{(N)}+\delta V^{(N)}\bigr{)}\widehat{a}^{(N)}=f^{(N)},$ (27)
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+u\cdot\gamma_{N}\Lambda_{N}{\lVert\widehat{a}^{(N)}\rVert}_{2},$ (29)
where $\Lambda_{N}$ denotes the Lebesgue constant for $Z$. If, in addition,
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+2u\cdot\gamma_{N}\Lambda_{N}{\lVert a^{(N)}\rVert}_{2}.$ (32)
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+{\lVert\widehat{P}_{N}-P_{N}\rVert}_{L^{\infty}(\Gamma)}$
$\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+\Lambda_{N}{\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}$
$\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+u\cdot\gamma_{N}\Lambda_{N}{\lVert\widehat{a}^{(N)}\rVert}_{2}.$ (33)
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+2u\cdot\gamma_{N}\Lambda_{N}{\lVert a^{(N)}\rVert}_{2}.$ (35)

### v5

@@
In this paper, we show that the monomial basis is generally as good as a well-conditioned polynomial basis for interpolation, provided that the condition number of the Vandermonde matrix is smaller than the reciprocal of machine epsilon. This leads to a practical algorithm for piecewise polynomial interpolation over general regions in the complex plane using the monomial basis. Our analysis also yields a new upper bound for the condition number of an arbitrary Vandermonde matrix, which generalizes several previous results.
@@ 1 Introduction
Function approximation has been a central topic in numerical analysis since its inception. One of the most effective methods for approximating a function $F:[-1,1]\to\mathbbm{R}$ is the use of an interpolating polynomial $P_{N}$ of degree $N$ which satisfies $P_{N}(x_{j})=F(x_{j})$ for a set of $(N+1)$ interpolation points $\{x_{j}\}_{j=0,1,\dots,N}$. In practice, the interpolation points are typically chosen to be the Chebyshev points, and the resulting interpolating polynomial, known as the Chebyshev interpolant, is a nearly optimal approximation to $F$ in the space of polynomials of degree at most $N$ [32]. A common basis for representing the interpolating polynomial $P_{N}$ is the Lagrange polynomial basis, and the evaluation of $P_{N}$ in this basis can be done stably using the Barycentric interpolation formula [9, 22]. Some other commonly used bases are Newton polynomials, Chebyshev polynomials, and Legendre polynomials. Alternatively, the monomial basis can be used to represent $P_{N}$, such that $P_{N}(x)=\sum_{k=0}^{N}a_{k}x^{k}$ for some coefficients $\{a_{k}\}_{k=0,1,\dots,N}$. The computation of the monomial coefficient vector $a^{(N)}:=(a_{0},a_{1},\dots,a_{N})^{T}\in\mathbbm{R}^{N+1}$ of the interpolating polynomial $P_{N}$ requires the solution of the linear system $V^{(N)}a^{(N)}=f^{(N)}$, where
$V^{(N)}:=\begin{pmatrix}1&x_{0}&x_{0}^{2}&\cdots&x_{0}^{N}\\ 1&x_{1}&x_{1}^{2}&\cdots&x_{1}^{N}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&x_{N}&x_{N}^{2}&\cdots&x_{N}^{N}\\ \end{pmatrix}\in\mathbbm{R}^{(N+1)\times(N+1)}$ (1)
@@
is a Vandermonde matrix, and $f^{(N)}:=\bigl{(}F(x_{0}),F(x_{1}),\dots,F(x_{N})\bigr{)}^{T}\in\mathbbm{R}^{N+1}$ is a vector of the function values of $F$ at the $(N+1)$ interpolation points on the interval $[-1,1]$. It is well-known that, given any set of real interpolation points within the unit interval, the condition number of $V^{(N)}$ grows at least as fast as $\sqrt{2}(1+\sqrt{2})^{N-1}/\sqrt{N+1}$ [8]. It follows that the accuracy of the numerical solution to this linear system deteriorates rapidly as $N$ increases, and, as a result, this algorithm for constructing $P_{N}$ is often considered unstable. But, is this really the case? Let $\{x_{j}\}_{j=0,1,\dots,N}$ be the set of $(N+1)$ Chebyshev points on the interval $[-1,1]$, and suppose that $F(x)=\cos(2x+1)$. We solve the resulting Vandermonde system using LU factorization with partial pivoting. In Figure 1(a), we show the polynomial interpolation error ${\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}$ and the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}$, where $P_{N}$ is approximated using the Barycentric interpolation formula and $\widehat{P}_{N}$ is evaluated using the computed monomial expansion. The Chebyshev interpolant constructed from the computed monomial expansion is, surprisingly, as accurate as the Chebyshev interpolant evaluated using the Barycentric interpolation formula (which is accurate up to machine precision), despite the huge condition number of the Vandermonde matrix reported in Figure 1(b).
(a) $F(x)=\cos(2x+1)$
(b) Condition number of $V^{(N)}$
@@
Figure 1: Polynomial interpolation of $\cos(2x+1)$ in the monomial basis over $[-1,1]$. The $x$-axis label $N$ denotes the order of approximation. The polynomial $P_{N}$ denotes the interpolating polynomial approximated using the Barycentric interpolation formula. The polynomial $\widehat{P}_{N}$ denotes the computed monomial expansion. The $L^{\infty}$ error is estimated by comparing the approximated function values at $10000$ equidistant points over $[-1,1]$ with the true function values.
@@
What happens when the function $F$ requires a higher-order polynomial for accurate approximation? In Figure 2, we compare the accuracy of the two approximations when $F(x)=\cos(8x+1)$ and when $F(x)=\cos(12x+1)$. Initially, the computed monomial expansion is as accurate as the Chebyshev interpolant evaluated using the Barycentric interpolation formula. However, the convergence of polynomial interpolation in the monomial basis stagnates after reaching a certain error threshold. It appears that, the higher the polynomial order necessary to approximate the functions, the larger that error threshold becomes. But consider the accuracy of the two approximations when $F(x)=\frac{1}{x-\sqrt{2}}$ and when $F(x)=\frac{1}{x-0.5i}$, shown in Figure 3. These two functions each have a singularity in a neighborhood of the interval $[-1,1]$, and Chebyshev interpolants of degree $\geq 40$ are required to approximate them to machine precision. Yet, no stagnation of convergence is observed. In Figure 4, we consider the case where $F$ is a non-smooth function, and we find that the accuracy of the two approximations is, once again, the same. The stability of polynomial interpolation in the monomial basis clearly depends in a subtle way on the function being approximated.
(a) $F(x)=\cos(8x+1)$
(b) $F(x)=\cos(12x+1)$
@@
Figure 2: Polynomial interpolation of more complicated functions in the monomial basis, over $[-1,1]$.
(a) $F(x)=\frac{1}{x-\sqrt{2}}$
(b) $F(x)=\frac{1}{x-0.5i}$
@@
Figure 3: Polynomial interpolation of functions with a singularity near the approximation domain, in the monomial basis, over $[-1,1]$.
(a) $F(x)={\lvert x+0.1\rvert}^{2.5}$
(b) $F(x)={\lvert\sin(5x)\rvert}^{3}$
@@
Figure 4: Polynomial interpolation of non-smooth functions in the monomial basis, over $[-1,1]$.
@@
These seemingly mysterious experiments can be explained partially from the point of view of backward error analysis. The forward error ${\lVert a^{(N)}-\widehat{a}^{(N)}\rVert}_{2}$ of the numerical solution $\widehat{a}^{(N)}$ to the Vandermonde system $V^{(N)}a^{(N)}=f^{(N)}$ can indeed be huge, but it is the backward error, i.e., ${\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}$, that matters for the accuracy of the approximation. This is because a backward error of size around machine precision implies that the difference between the computed monomial expansion, which we denote by $\widehat{P}_{N}$, and the exact interpolating polynomial, $P_{N}$, is a polynomial which approximately vanishes at all of the interpolation points. More specifically, the residual error ${\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}$ is bounded by $\Lambda_{N}\cdot{\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}$, where $\Lambda_{N}$ denotes the Lebesgue constant associated with the interpolation points. This constant is defined by the formula $\Lambda_{N}=\sup_{F\in C([-1,1]),F\neq 0}\frac{{\lVert P_{N}\rVert}_{\infty}}{{\lVert F\rVert}_{\infty}}$, with ${\lVert\cdot\rVert}_{\infty}$ denoting the $\infty$-norm in $C([-1,1])$, and $P_{N}$ denoting the interpolating polynomial of $F$ for the given interpolation points. This constant grows logarithmically when well-conditioned interpolation points, such as Chebyshev points, are used. As a result, we can bound the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}$ using the following inequality:
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}+{\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}([-1,1])}$
$\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}+\Lambda_{N}\cdot{\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}$
$\lesssim$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}([-1,1])}+{\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}.$ (2)
Here, we introduce the notation $x_{N}\lesssim y_{N}$ to indicate the existence of a constant $c_{N}>0$, independent of the specific values of $x_{N}$ and $y_{N}$ but dependent on $N$, such that $x_{N}\leq c_{N}y_{N}$. This constant $c_{N}$ is assumed to be a rational function of $N$, where $N$ represents either the dimensionality or the approximation order. Additionally, we write $x_{N}\approx y_{N}$ to mean that both $x_{N}\lesssim y_{N}$ and $y_{N}\lesssim x_{N}$.
@@
When the backward error is smaller than the polynomial interpolation error, the monomial approximation error is dominated by the polynomial interpolation error, and the use of a monomial basis does not incur any additional loss of accuracy. Once the polynomial interpolation error becomes smaller than the backward error, the convergence of the approximation stagnates. For example, for the function approximated in Figure 3(a), we find that the backward error is around the size of machine epsilon for all $N\leq 43$, which explains why no stagnation is observed. On the other hand, in Figure 2(a), the backward error is around the size of $10^{-13}$ for $N\geq 20$, which leads to stagnation once the polynomial interpolation error is less than $10^{-13}$.
@@
The explanation above brings up a new question: when will the backward error be small? Given a matrix $A\in\mathbbm{C}^{n\times n}$ and a vector $b\in\mathbbm{C}^{n}$, the numerical solution $\widehat{x}$ to the linear system $Ax=b$ computed by a backward stable solver is the exact solution to the linear system
$(A+\delta A)\widehat{x}=b$ (3)
for a perturbation matrix $\delta A\in\mathbbm{C}^{n\times n}$ that satisfies ${\lVert\delta A\rVert}_{2}\leq u\cdot\zeta_{n}{\lVert A\rVert}_{2}$, where $u$ denotes machine epsilon and $\zeta_{n}>0$ is a small growth factor that is independent of $A$ and $b$ (see Section 7.6 in [21]). A classical example of a backward stable linear system solver is the Householder QR factorization, whose growth factor $\zeta_{n}$ is bounded by a low-degree polynomial in $n$ (see Theorem 19.5 and Equation (19.14) in [21]). Another commonly used solver, LU factorization with partial pivoting, has a growth factor $\zeta_{n}$ that can grow exponentially as $n$ increases (see Section 9.3 in [21]). This upper bound for ${\lVert\delta A\rVert}_{2}$ is, however, rarely approached, and the stability of LU factorization with partial pivoting is often comparable to that of QR factorization in practice (see Lecture 22 in [34] for a detailed discussion). When we state in this paper that a linear system is solved using a backward stable solver, we assume that the growth factor $\zeta_{n}$ associated with this solver is bounded by a low-degree polynomial in $n$ with small coefficients.
When the Vandermonde system $V^{(N)}a^{(N)}=f^{(N)}$ is solved using a backward stable solver, the backward error of the numerical solution to the Vandermonde system satisfies ${\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}={\lVert\delta V^{(N)}\widehat{a}^{(N)}\rVert}_{2}\leq u\cdot\zeta_{n}{\lVert V^{(N)}\rVert}_{2}{\lVert\widehat{a}^{(N)}\rVert}_{2}$. Since ${\lVert V^{(N)}\rVert}_{2}\in[\sqrt{N+1},N+1]$ when the interpolation points are inside the unit interval $[-1,1]$, the backward error is essentially dominated by $u\cdot{\lVert\widehat{a}^{(N)}\rVert}_{2}$. Furthermore, one can show that ${\lVert\widehat{a}^{(N)}\rVert}_{2}\approx{\lVert a^{(N)}\rVert}_{2}$ when the condition number of $V^{(N)}$ satisfies $\kappa(V^{(N)})\lesssim\frac{1}{u}$ (see Theorem 2.2 for a formal statement). It follows that the monomial approximation error can be quantified a priori using information about the interpolating polynomial, which means that a theory of polynomial interpolation in the monomial basis can be developed. Although the examples provided in this section are limited to the interval $[-1,1]$, it is important to note that these observations are equally applicable to a simply connected compact domain in the complex plane. In the rest of the paper, we consider polynomial interpolation in the monomial basis over a simply connected compact domain in the complex plane, with the interval as a special case.
@@
The rest of the paper is organized as follows. In Section 2, we analyze polynomial interpolation in the monomial basis over a simply connected compact domain in the complex plane. Our analysis reveals the conditions under which the monomial basis is as good as a well-conditioned polynomial basis for interpolation, resulting in a practical algorithm for using the monomial basis, with no extra error and with almost no extra cost. As a by-product of our analysis, we derive a tight upper bound for the condition number of any Vandermonde matrix (see Theorem 2.5). In Section 3, we provide extensive numerical experiments to support our analysis. In Section 4, we present applications where the use of a monomial basis for interpolation offers a substantial advantage over other bases. In Section 5, we summarize our key points, review related work, and discuss generalizations of our results.
@@ 2 Polynomial interpolation in the monomial basis
In this section, we consider polynomial interpolation of a function over a simply connected compact set in the complex plane. We structure the section as follows. First, we show in Theorem 2.2 that, when $\kappa(V^{(N)})\lesssim\frac{1}{u}$, the monomial approximation error is bounded by the sum of the polynomial interpolation error and a extra error term that involves the 2-norm of the monomial coefficient vector of the interpolant. Then, we provide tight upper bounds for this extra error term and for the growth of the condition number of the Vandermonde matrix in Theorems 2.4 and 2.5, respectively. In Section 2.1, we elucidate the frequently observed and seemingly paradoxical success of polynomial interpolation in the monomial basis when $\kappa(V^{(N)})\lesssim\frac{1}{u}$. Finally, we summarize the necessary considerations for the proper usage of the monomial basis in Section 2.2.
Let $\Omega\subset\mathbbm{C}$ be simply connected and compact, and let $F:\Omega\to\mathbbm{C}$ be an arbitrary function. The $N$th degree interpolating polynomial, denoted by $P_{N}$, of the function $F$ for a given set of $(N+1)$ distinct interpolation points $Z:=\{z_{j}\}_{j=0,1,\dots,N}\subset\Omega$ can be expressed as $P_{N}(z)=\sum_{k=0}^{N}a_{k}z^{k}$, where the monomial coefficient vector $(a_{0},a_{1},\dots,a_{N})^{T}$ is the solution to the Vandermonde system
@@
In this paper, we choose to fix the set of basis functions as $\{z^{k}\}_{k=0,1,\dots,N}$. However, depending on the geometry and location of $\Omega$, it can be advantageous to use a scaled and translated monomial basis $\{(\frac{z-\mu}{\nu})^{k}\}_{k=0,1,\dots,N}$. Our analysis also encompasses this case, as the Vandermonde system corresponding to the scaled and translated basis with the points $\{z_{j}\}_{j=0,1,\dots,N}\subset\Omega$ is equivalent to that of the original basis $\{z^{k}\}_{k=0,1,\dots,N}$ using the transformed points $\{\frac{z_{j}-\mu}{\nu}\}_{j=0,1,\dots,N}\subset\frac{\Omega-\mu}{\nu}$. We assume without loss of generality that $\Omega$ lies within the unit disk, ensuring that ${\lVert V^{(N)}\rVert}_{2}\in[\sqrt{N+1},N+1]$.
In the following, we denote the Vandermonde matrix, the monomial coefficient vector, and the right-hand side vector by $V^{(N)}$, $a^{(N)}$, and $f^{(N)}$, respectively.
@@
${\lVert\delta A\rVert}_{2}\leq\frac{1}{\alpha\cdot{\lVert A^{-1}\rVert}_{2}},$ (5)
@@
The proof is completed by combining (9) and (10). $\blacksquare$
The following theorem provides upper bounds for the monomial approximation error. It is related to frame approximation [3, 4] and the method of fundamental solutions [6, 31] in the sense that it also uses a similar idea, that it is possible to achieve a small approximation error with an ill-conditioned basis, provided that there exists a small coefficient vector with a good fit.
@@
Suppose that there exists some constant $\gamma_{N}\geq 0$ such that the computed monomial coefficient vector $\widehat{a}^{(N)}=(\widehat{a}_{0},\widehat{a}_{1},\dots,\widehat{a}_{N})^{T}$ satisfies
$\bigl{(}V^{(N)}+\delta V^{(N)}\bigr{)}\widehat{a}^{(N)}=f^{(N)}$ (11)
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+u\cdot\gamma_{N}\Lambda_{N}{\lVert\widehat{a}^{(N)}\rVert}_{2},$ (13)
where $\Lambda_{N}:=\sup_{F\in C(\Omega),F\neq 0}\frac{{\lVert P_{N}\rVert}_{L^{\infty}(\Omega)}}{{\lVert F\rVert}_{L^{\infty}(\Omega)}}$ is the Lebesgue constant for $Z$. If, in addition,
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+2u\cdot\gamma_{N}\Lambda_{N}{\lVert a^{(N)}\rVert}_{2}.$ (16)
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+{\lVert\widehat{P}_{N}-P_{N}\rVert}_{L^{\infty}(\Omega)}$
$\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+\Lambda_{N}{\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}$
$\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+u\cdot\gamma_{N}\Lambda_{N}{\lVert\widehat{a}^{(N)}\rVert}_{2}.$ (17)
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\leq$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+2u\cdot\gamma_{N}\Lambda_{N}{\lVert a^{(N)}\rVert}_{2}.$ (19)